Anna, Ben, and Cindy have M markers together. Anna has A markers, and this is twice as many markers as Ben has. How many markers does Cindy have?
_______markers
Translate into Algebra.

Answer by stanbon(75887)

Anna, Ben, and Cindy have M markers together. Anna has A markers, and this is twice as many markers as Ben has. How many markers does Cindy have?
_______markers
Translate into Algebra.
------------
Equations:
A + B + C = M
A = a
A = 2B
-----
2B + B + C = M
----
Ans:: C = M - 3B
=============
Cheers,
Stan H.
=============